OpenStudy (anonymous):

where does the following function intersect the x-axis. f(x)=-9x^2-9x-1.

OpenStudy (anonymous):

\[\left\{x\to \frac{1}{6} \left(-3-\sqrt{5}\right),x\to \frac{1}{6} \left(-3+\sqrt{5}\right)\right\}=\{x\to -0.872678,x\to -0.127322\} \]Visual aid attached.